For the two events $A$ and $B$,$P(A) = 0.38$ and $P(B) = 0.41$. What is the value of $P(\text{not } A)$?

  • A
    $0.41$
  • B
    $0.62$
  • C
    $0.59$
  • D
    $0.21$

Explore More

Similar Questions

Two numbers are selected randomly from the set $S = \{ 1, 2, 3, 4, 5, 6 \}$ without replacement one by one. The probability that the minimum of the two numbers is less than $4$ is

An urn contains $25$ balls numbered $1$ to $25$. Suppose an odd number is considered a 'success'. If $2$ balls are drawn from the urn with replacement,find the probability of getting at most $2$ successes.

$A$ box contains $4$ white pens and $2$ black pens. Another box contains $3$ white pens and $5$ black pens. If one pen is selected from each box,then the probability that both the pens are white is equal to

In a single throw of two dice,what is the probability of obtaining a sum of numbers greater than $7$,given that $4$ appears on the first die?

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo